Solve each of the following systems by using either addition or substitution. If a unique solution does not exist, state whether the system is dependent or inconsistent.

10x + 2y = 7
y= -5x + 3

Answer by scott8148(6628)

multiplying the second equation by -2 gives -2y=10x-6 ... adding this result to the first equation gives

10x+2y+(-2y)=7+(10x-6) ... 10x=10x+1 ... this is an inconsistent result